OpenStudy (anonymous):

Consider the functions f( x) = square root of x and g(x). =7x+ b. In the standard (x,y) coordinate plane, y = f( g ( x)) passes through (4,6). What is the value of b?

OpenStudy (daniel.ohearn1):

So how would you first go about writing f(g(x))? Any ideas?

OpenStudy (kkutie7):

\[f(x)=\sqrt{x}\] \[g(x)=7x+b\] it looks like you need to find the new function y

OpenStudy (kkutie7):

same concept as this: \[f(x)=x^{2}\] \[x=2\] and we need y=f(x) \[y=f(x)=2^{2}=4\]

OpenStudy (kkutie7):

you need to substitute g(x) into the function f(x). from there you need to use the coordinates to find your unknown. once you have : \[y=f(g(x))\] you can do this: \[6=f(g(4))\] tag me if you need help, but I encourage you to type what you can do in here first =)

jhonyy9 (jhonyy9):

f(x)=sqrt(x) g(x)=7x+b f(g(x)) = sqrt(g(x)) = sqrt(7x+b) from y=f(g(x)) pass. trough (4,6) mean that f(g(4))=6 so sqrt(7*4+b) = 6 sqrt(28+b) = 6 square both sides 28+b=36 b=36-28 b=8 hope this will help you so easy understanding the way hhow you can solve it right